OpenStudy (anonymous):

suppose cos=1/2 and sin x<0. What is the value of tan x?

OpenStudy (perl):

since cos is positive and sin is negative, you are in quadrant 4

OpenStudy (anonymous):

if \(\cos(x)=\frac{1}{2}\) and sine is negative, then \(x=\frac{5\pi}{3}\)

OpenStudy (ajspeller):

Cosine is positive and Sine is negative in Quadrant IV.